Determine the value of the equilibrium constant, Q for the chemical reaction given below, if 0.230 moles of Mg, 0.087, sulfate (SO4) 0.860, CO2 with 0.0580 and 0.120 moles of water vapor are there in a 5 liters container and at an equilibrium.


Mg= 0.230


SO4= 0.860


CO2= 0.0580


water vapor or H2O= 0.120


container 5 liters


                              


substitute the value of the compounds


1
Expert's answer
2021-04-05T07:49:31-0400

"MgCO_3 + H_2SO_4 = MgSO_4 + CO_2 + H_2O"


Molar Conc. of Mg = 0.087/5

Molar Conc. of MgCO3 =

Molar Conc. of SO4 = 0.230/5

Molar Conc. of H2SO4 = 0.230/5


Molar Conc. of CO2 = 0.0580/5

Molar Conc. of H2O= 0.230/5


"Q = \dfrac{0.875/5×0.230/5}{0.0580/5×0.230/5}= 15.1"